OpenStudy (sleepyjess):

Proving trig equations \(\sf 1+sec^2xsin^2x=sec^2x\)

OpenStudy (sleepyjess):

I am thinking to use \(\sf 1+tan^2x=sec^x\) but I don't think that is right.

OpenStudy (sleepyjess):

That should be \(\sf sec^2x\)

ganeshie8 (ganeshie8):

You're on right track.. keep going :)

ganeshie8 (ganeshie8):

Take the left hand side \[1+\sec^2x\sin^2x\] To use the identity, somehow you need to make the second term \(\sec^2x\sin^2x\) equal to \(\tan^2 x\). Maybe start by writing sec^2x as 1/cos^2x

OpenStudy (sleepyjess):

Ok, \(\sf 1+\dfrac{1}{cos^2x}sin^2x=sec^2x\). I believe I somehow have to make \(\sf \dfrac{1}{cos^2x}sin^2x\) become \(\sf\dfrac{sin^2x}{cos^2x}\)

ganeshie8 (ganeshie8):

to multiply two fractions, you simply multiply the numerators and denominators separately : \[\dfrac{1}{a}\cdot b = \dfrac{1}{a}\cdot \dfrac{b}{1} = \dfrac{1\cdot b}{a\cdot 1} = \dfrac{b}{a}\]

ganeshie8 (ganeshie8):

In short \[\dfrac{1}{a}\cdot b = \dfrac{b}{a}\]

OpenStudy (solomonzelman):

\(\large\color{black}{1+\sec^2x \sin^2x=\sec^2x}\) \(\large\color{black}{\cos^2x\sec^2x+\sec^2x \sin^2x=\sec^2x}\) \(\large\color{black}{(\cos^2x+ \sin^2x)\sec^2x=\sec^2x}\)

OpenStudy (solomonzelman):

like this

OpenStudy (solomonzelman):

ganesh, what do you think?:P

ganeshie8 (ganeshie8):

thats kinda cute :D

OpenStudy (solomonzelman):

:D

OpenStudy (sleepyjess):

Oh, I forgot that. So after multiplying there is \(\sf 1+\dfrac{sin^2x}{cos^2x}=sec^2x\). \(\sf\dfrac{sin^2x}{cos^2x}=tan^2x\) Then \(\sf 1+tan^2x=sec^2x\) O_O

OpenStudy (loser66):

yuuuuuup, so simple, right?

OpenStudy (solomonzelman):

oh, just rewriting it as tan^2x is definitely better

OpenStudy (solomonzelman):

LoL

OpenStudy (solomonzelman):

I haven't recognized that it is tan^2x sitting in there...